OpenStudy (anonymous):

What is the approximate length of the base of an isosceles triangle if the congruent sides are 2.5 m and the vertex angle is 25°?

OpenStudy (anonymous):

law of cosines:\[c^2=a^2+b^2-2ab~cos(C)\]

OpenStudy (anonymous):

in this case a=b\[c^2 = 2a^2-2a^2~cos(C)\] in this case a=b\[c^2 = 2a^2(1-cos(C))\]